To determine whether a shiny gold-colored rock is actually gold, a chemistry student decides to measure its heat capacity. She first weighs the rock and finds it has a mass of 4.7 g. She then finds that upon absorption of 52.7 J of heat, the temperature of the rock rises from 25 ∘C to 57 ∘C.

Find the specific heat capacity of the substance composing the rock. Express the specific heat capacity in joules per gram-Celsius to two decimal places.

Complete previous part(s) Part C A 55.0 g aluminum block initially at 27.5 ∘C absorbs 725 J of heat.

What is the final temperature of the aluminum? Express your answer in degrees Celsius to one decimal place.
Answer #1

we know the formula

q = mass of substance   specific heat of substance      T

where,

q = heat = 52.7 J

T = change in temperature =  57 - 25 = 32 0C

specific heat of rock = ?

mass of rock = 4.7 gm

we can write above formula

specific heat of rock = q / ( mass of rock ) X ( T)

substitute the value

specific heat of rock = 52.7 J / ( 4.7 gm X 32 0C ) = 0.35 J / g 0C

specific heat of rock = 0.35 J / g 0C

part C

we know the formula

q = mass of substance   specific heat of substance      T

where,

q = heat = 725 J

T = change in temperature = ?

specific heat of Aluminium = 0.902 J/g0C

mass of Aluminium = 55 gm

Initial temperature = 27.5 0C

we can write above formula

T = q / (mass of aluminium X specific heat of aluminium)

substitute the value

T = 725 J / ( 55 g X 0.902 J / g 0C ) = 14.6 0C

final temperature = initial temperature + T  

final temperature of aluminium = 27.5 + 14.6 = 42.1 0C

final temperature of aluminium = 42.1 0​​​​​​​C